A 50-turn circular coil (radius = 15 cm) with a to

please show calculations. the answers are 11mA, 4.8uC, and 1.0V
Answer #1

1)

emf=d?/dt

current i=emf/R

R=4

?=N*pi*r^2*Asin(alpha*t)

N=50 r=0.15

A=80*10^-6 alpha=50 pi

d?/dt=N*pi*r^2*A*alpha*cos(alpha*t)

at t=20ms

d?/dt=50*pi*0.15*0.15*80*10^-6**50*pi*-1

=44.4132198049*10^-3


i=44.4132198049*10^-3/4

=11 mA


2)

B due to long wire at distance x is

B=